A sample of sodium percarbonate (Na2CO3*nH2O2) weighing 0.024g is diluted using a 0.10 M solution of sodium thiosulfate (Na2S2O3). The volume of sodium thiosulfate used is 4.25mL. There are 3 questions:

1. What is the amount of hydrogen peroxide in the sample in moles?

2. What is the amount of sodium carbonate in the sample in moles?

3. What is the molar ratio of H2O2:Na2CO3?

So far, I've found that there are 0.000425 mol Na2S2O3 ; however, I'm unsure how to back out the amount of hydrogen peroxide and sodium carbonate?
Answer #1

the chemical formula of sodium percarbonate is 2Na2CO3.3H2O2. 2 moles of Sodium carbonate crystallise with 3 moles hydrogen peroxide forming an adduct.

the molecular weight of the adduct is 2*[2*22.99 + 12.011 + 3*15.999] + 3*[2*1.008 + 2*15.999] = 211.976 + 102.042 = 314.396 gm/mole.

1. Fraction of H2O2 in the adduct = 102.042/314.396 = 0.3258

Amount of H2O2 present in 0.024gm of the adduct = 0.3258*0.024 = 0.0078 gm

Molecular weight of H2O2 is [2*1.008 + 2*15.99] = 34.014 gm/mole

No. of moles of H2O2 present in the provided sample = 0.0078/34.014 = 2.29*10-4 moles

2. Fraction of Na2CO3 in the adduct = 211.976/314.396 = 0.6742

Amount of Na2CO3 present in 0.024gm of the adduct = 0.6742*0.024 = 0.0162 gm

Molecular weight of Na2CO3 is [2*22.99 + 12.011 +3*15.999] = 105.988 gm/mole

No. of moles of Na2CO3 present in the provided sample = 0.0162/105.988 = 1.53*10-4 moles

3. Molar ratio of H2O2:Na2CO3 = 2.29*10-4/1.53*10-4 = 1.5496